Circular No. 8450

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2004 X2 (LINEAR) An apparently asteroidal object discovered by the LINEAR project

(discovery observation below), and posted on the ‘NEO Confirmation Page’, has been found to be cometary in appearance. R-band observations taken by C. Hergenrother with the Catalina 1.54-m reflector on Dec. 9.55 UT show a coma 25′′ in diameter and a slightly curved 50′′ tail in p.a. 325^{o}. CCD images taken by J. E. McGaha (Tucson, AZ, 0.36-m f/10 reflector) on Dec. 10.4 show a 3′′ coma with a 12′′ tail in p.a. 300^{o}.

2004 UT  α_{2000}  δ_{2000}  Mag.
Dec. 8.46681  9^{h}31^{m}07.^{s}95  +12^{o}34′01′′.7  18.9

The available astrometry (including LONEOS observations from Nov. 19), the following parabolic orbital elements, and an ephemeris appear on MPEC 2004-X35.

T = 2004 Aug. 25.890 TT  ω = 162.^{o}526
Ω = 307.353 }2000.0
q = 3.79930 AU  i = 72.173

SUPERNOVA 2004gp A. Garg, M. Huber, P. Garnavich, P. Challis, C. Stubbs, and the Super-

MACHO collaboration (cf. IAUC 8439) report the detection with the Cerro Tololo 4-m telescope (+ MOSAIC imager) on Oct. 19 of a supernova (magnitude ‘VR’ ∼ 21.0 at peak) located at α = 4^{h}53^{m}38^{s}.96, δ = −68^{o}47′02′′.8 (equinox 2000.0), which is 0′′.28 east and 0′′.12 north of a galaxy (whose magnitude is I = 18.7) situated behind the Large Magellanic Cloud. The new object was also detected in 4-m-telescope observations on Sept. 18, Oct. 5, 11, 17, 23, and Nov. 6. Spectroscopy obtained on the Magellan Clay telescope (+ LDSS-2 spectrograph) on Nov. 6 indicates that SN 2004gp is a type-II supernova at a redshift of 0.067.

SUPERNOVA 2004go IN IC 270 N. Morrell, M. Hamuy, G. Folatelli, and M. Phillips, Carnegie Super-

nova Project, report that a CCD spectrogram (range 380–923 nm) of SN 2004go (cf. IAUC 8448), obtained on Dec. 9.23 UT with the Las Campanas du Pont 2.5-m telescope (+ WFCCD spectrograph), shows it to be a type-Ia supernova, probably 3 to 4 weeks after maximum light.

2004 December 10  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2004 CBAT  Daniel W. E. Green